29. The following table summarizes the number a minutes a certain bus is late:

         x

0

1

2

3

p(x)

0.62

0.24

0.11

0.03

The mean number of minutes the bus is late equals:

         [A]     0                [B]     1.5              [C]     0.95            [D]    0.55
Answer #1

Solution :

probability distribution

Mean = = X * P(X)

= 0 * 0.62 + 1 * 0.24 + 2 *0.11 + 3 * 0.03

= ( 0+ 0.24 + 0.22 + 0.09)

= 0.55